2.65 A car and a truck start from rest at the same instant, with the car initially at some distance behind the truck. The truck has a constant acceleration of 2.10 m/s, and the car has an acceleration of 3.40 m/s2. The car overtakes the truck after the truck has moved 60.0 m. (a) How much time does it take the car to overtake the truck? (b) How far was the car behind the truck initially? (c) What is the speed of each when they are abreast? (d) On a single graph sketch the position of each vehicle as a function of time. Take x 0 at the initial location of the truck.
media%2F639%2F639337fa-53fe-45a6-bfba-19
Answer #1

2.65

Time taken by truck to cover 60 m =sqrt(2*60/2.1)=7.6 sec

So car will overtake in 7.6 sec

(b) distance travelled by car=0.5*3.4*7.6^2=97.14 m

So car is 37.14 m behind truck

(c) speed of car=3.4*7.6=25.84 m/s

Speed of truck=2.1*7.6=15.96 m/s

Truck 6 Cat